# Representing Molecules as Random Walks Over Interpretable Grammars

### v1

@@ 4.4.2 Two-dimensional T-SNE on $H_{M}$ versus pretrained representations
To answer question 7), we analyze the 2D t-SNE embeddings of various methods’ final layer representations of 64 test set molecules on HOPV. As shown in Figure 7, our method is unique in extracting visually meaningful representations. High HOMO molecules were identified from the visual clusters for structural analysis. Molecules in the upper cluster as illustrated in Figure 8 often have structures promoting electron delocalization, like carbonohydrazonoyl dicyanide, while those in the lower cluster have electron-donating groups or structures increasing steric hindrance to boost HOMO values as shown in Figure 8. These two structural features correspond to the two primary ways to design molecules with high HOMO values. These findings aid the search for novel molecules with desirable photovoltaic properties.
@@ 4.5 Conclusion & Future Work
We represent molecules as random walks over an interpretable context-sensitive grammar over the motif graph, a hierarchical abstraction over the design space. Our evaluation on downstream property prediction and molecular generation tasks shows our representation combines quantitative advantages in performance and efficiency with qualitative advantages of simplicity and enhanced interpretability. We devise and execute a practical workflow that invites experts in the loop to enhance our design basis and representations by fragmenting molecules into well-established functional groups, creating a synergy between expert feedback and the quality of our representations. One promising avenue of future research is improving the autonomous extraction of motifs through learnable and/or human-guided approaches with Large Language Models.
5 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Expert Annotation Workflow
@@
First, experts view figures of the molecules, and indicate the bonds to break in order to segment the molecule into coherently chosen sub-fragments, shown in 11. We provide a brief description of the example datasets we show here and elaborate on the rationale behind the experts’ segmentation strategy:
Table 5: Segmentation of the molecules (a) to (d) in Figure 11. Bonds to break indicates the chemical bonds to cut to create black fragments, while the black groups and red groups listed for each molecules correspond to one another, respectively.
@@
Predictive Toxicology Challenge (PTC) (Helma et al., 2001). The small molecules are characterized by distinct functional groups known for their carcinogenic properties or liver toxicity (Miller et al., 1949; Helma et al., 2001). These groups comprise a rich variety of elements such as halides, alkylating agents, epoxides, and furan rings. (Figure 3) Therefore, we specifically segmented it into functional groups and sub-structures that contribute most to the toxicity of the compounds (Hughes et al., 2015).
@@ Appendix B Representing Existing Molecules as Walks on This Graph
Extracting Walks from Segmentation (HOPV, PTC)
During segmentation, we use the rules in step 2 of Appendix section (A) to segment a molecule into fragments. In doing so, we also obtain the molecule’s representation $H_{M}$ as a directed subgraph over the motif graph. The pseudocode is found in Algorithm 1.
@@ 13 Q $\leftarrow$ queue([(root,src)]);
Algorithm 2 linearizes the molecule into a directed acyclic graph (DAG). This procedure begins by finding the longest path, and choosing a consistent ordering over neighbors ($N_{G_{i}}$) to determine the random walk sequence. We elaborate on the reasoning behind this canonicalization in Appendix B. The DAG constraint enables our graph diffusion process to become a generator of new molecules (as will be discussed in Appendix D), in addition to capturing the distribution of existing ones. Thus, we specifically ask experts to create segmentations that are acyclic, which they naturally do in nearly all cases anyway. In the case of monomers, this canonicalization is consistent with the IUPAC nomenclature(IUPAC, 1997) of linearizing a monomer via its longest (main) chain, where $N_{G_{i}}$ should iterate over neighbor fragments that descend side chains before the consecutive fragment on the backbone of the main chain. More specifically, src and dest in Algorithm 2 correspond to the first and last group of the main chain.
Extracting Walks From Literature (Case Study of Group Contribution)
@@
The Group Contribution dataset includes a compilation of motifs characterized for gas separation, including common organic chemical functional groups as well as important scaffold functional groups such as Triptycene and its derivatives, dioxin and its derivatives, and N-methylphthalimide and PIM-1 and its derivatives (Figure 3) (Wang et al., 2018, 2020). These functional groups contribute significantly to maintaining the structures and properties of 3D scaffold building blocks in polymer self-assembly, which in turn play a significant role in gas separation processes, i.e. the separation of $H_{2},H_{2}/N_{2},O_{2},O_{2}/N_{2},CO_{2},CO_{2}/CH_{4}$ which are common separation tasks important in gas and oil industry. The steps we take for compiling this dataset of segmented monomers are as follows:
@@
Graph Augmentation
The motif graph is the directed, multi-edge graph $G=(V,E)$. When traversing to a previously seen motif v, there is ambiguity in whether the random walk forms a cycle vs creating a copy of the previous motif and appending to the trajectory. To remove this ambiguity, the random walk traverses a duplicate node, $v_{k}$ for the latter case. A dataset of molecules and their representations, $D:=\{(M,H_{M})\}$ thus induces an augmented version of $G=:G’$ ’. For each $v\in V$, let $K_{v}=max(count(v,H_{M})$ for M). We create duplicates of v: $V’:=V\cup\bigcup_{v\in V}{v_{k}\mid k=0,\ldots,k-1}$ $E’=E\cup\bigcup_{v\in V}{(v_{k},v’,e)\mid(v,v’,e)\text{ in }E,\forall k=0,\ldots,K_{v}-1}$ Molecule $M=(V_{M},E_{M})$ is then a rooted subgraph of $G’$ ’.
Data Augmentation
@@
As described in 2, we linearize a molecule by first setting its “main chain” – the longest shortest path of $H_{M}$. If this happens to be part of a cycle, we disregard one edge. If there are multiple longest shortest paths, we choose the one whose first differing node comes first in our canonical ordering over the nodes of G.
@@
Appendix C Building the motif graph
Expanding on Section 3.1, we apply a subgraph-matching algorithm with pseudocode in 3 over all pairs of motifs $v_{1},v_{2}$. This algorithm is embarrassingly parallel and runtime-efficient as the subgraph $v_{R_{l}}$ is, unless specified otherwise, a few atoms or a ring. RDKit(Landrum, 2016) provides out-of-the-box implementations for subgraph matching optimized for molecular sub-fragments like rings, enabling a significant speedup in runtime.
@@ 1 G = graph(V);
Figure 11: In Figure 3, $\{v_{1}\}_{r_{1}}=\{1,9\},\{v_{1}\}_{r_{2}}=\{1\},\{v_{1}\}_{r_{3}}=\{9\},\{v_{1}\}_{r_{4}}=\{5\}$. $\{v_{2}\}_{r_{1}}=\{1,2,3,4,5,6\},\{v_{2}\}_{r_{2}}=\{12,13,14,15,16,17\}$. We annotate $e_{1,1}$, where $b_{1}=\{6,4,3,2,8,7\},b_{2}=\{10,7\}$ provides the “certificate” of a successful match.
@@ C.2 Connection to Graph Coarsening
Mathematically, the motif graph advocated in this work is the quotient graph of the molecular graph, under the equivalence relation defined as $u\equiv v$ if nodes $u$ and $v$ belong to the same motif. As our motifs do not overlap and they jointly cover all nodes of the molecular graph, they define a partitioning of the graph. In scientific computing, collapsing each partition into a single node and retaining edges crossing partitions is called graph coarsening, which is a commonly used technique to solve large-scale problems, notably solving sparse linear systems of equations (Chen et al., 2022). Working on the coarsened version of the graph (i.e., the quotient graph) is computationally attractive as the graph size is much smaller. Moreover, when applied to machine learning problems such as graph classification, it is demonstrated that the representation learned from the quotient graph can be as predictive as that learned from the original graph (Chen et al., 2023; Ma & Chen, 2021; Cai et al., 2021). Favorably, a unique scenario of this work is that all concerned (molecular) graphs share the same set of motifs, which brings in the potential benefit of learning better molecule representations based on motif representations that form the basis of all molecules.
@@ Appendix D Grammar Learning
Graph Diffusion Strategy
@@ 6       $c^{(0)}\leftarrow$ [0 for v in G];
13             $c^{(t+1)}=\frac{t}{t+1}\cdot(c^{t}+\frac{1}{t+1})\cdot p^{t}$;
@@ 24             W $\leftarrow$ W - $\alpha*\frac{\text{dLoss}}{\text{dE}}$;
Visualizing Learning Process
@@
Figure 13: We show the weight evolution of the edges incidental to $G14$ on HOPV. (Left) After processing the raw dataset into random walks, we visualize the empirical distribution of edge traversals. (Middle) After learning our context-sensitive grammar, we plot the prior edge weights, i.e. the context-free parameter $E$. (Right) We plot the transition probabilities starting at $G14$ during the random walk generation process.
From Figure 12 and Figure 13, we see our grammar’s capacity to estimate the prior edge weights, $E$, through training, as well as correct the edge weights via a context-sensitive adjustment during the random walk. Weights of edges that are commonly traversed after $G14$ will be amplified during training, and weights of edges that visit $G14$ from another state will be diminished.
@@ Appendix E Property Prediction
Graph Neural Network Design Choices
Figure 14: Default hyperparameter settings for property prediction
@@
We apply a Graph Isomorphism Network (Xu et al., 2019) with hyperparameters in Table 14. For molecule M with representation $H_{M}$, the node-level features include: a) the Morgan fingerprint of the motif $v_{i}$ (dimension 2048), b) the context-free weights of its out-edges (dimension $|G|$), i.e. E[i]. We also concatenate the Morgan fingerprint of M.
Bag-of-Motifs Design Choices
We obtain the $|G|$ motif-occurrence feature vector for each molecule. Similar to our method, we concatenate the Morgan fingerprint of $M$ to it. We use XGBoost’s implementation with $16$ estimators (boosting rounds) and a maximum tree depth of $10$.
Optimization Design Choices

### v2

@@ 4.4.2 Two-dimensional T-SNE on $H_{M}$ versus pretrained representations
To answer question 7), we analyze the 2D t-SNE embeddings of various methods’ final layer representations of 64 test set molecules on HOPV. As shown in Figure 7, our method is unique in extracting visually meaningful representations. High HOMO molecules were identified from the visual clusters for structural analysis. Molecules in the upper cluster as illustrated in Figure 8 often have structures promoting electron delocalization, like carbonohydrazonoyl dicyanide, while those in the lower cluster have electron-donating groups or structures increasing steric hindrance to boost HOMO values as shown in Figure 8. These two structural features correspond to the two primary ways to design molecules with high HOMO values. These findings aid the search for novel molecules with desirable photovoltaic properties. As 2D t-SNE is not a universal way to analyze representations, we also visualize the agreement between embedding similarity and structural similarity using a $64\times 64$ grid. This is can be found in Appendix G.3, as part of an in-depth case study on HOPV.
@@ 4.5 Conclusion & Future Work
We represent molecules as random walks over an interpretable context-sensitive grammar over the motif graph, a hierarchical abstraction over the design space. We devise and execute a practical workflow that invites experts in the loop to enhance our design basis and representations by fragmenting molecules into well-established functional groups, creating a synergy between expert feedback and the quality of our representations. Our evaluation on downstream property prediction and molecular generation tasks shows our representation combines quantitative advantages in performance and efficiency with qualitative advantages of simplicity and enhanced interpretability. One promising avenue of future research is improving the autonomous extraction of interpretable grammar rules through learnable and/or human-guided approaches with Large Language Models.
Acknowledgements
This work is supported by the MIT-IBM Watson AI Lab and its member companies Evonik and Shell.
Impact Statement
This paper presents work whose goal is to concurrently and conjointly advance the fields of Machine Learning and Chemical Discovery. The application of our method can have consequences for real-world discovery workflows. There are no ethical aspects which we foresee and feel must be discussed here.
Appendix A Motif Collection Strategy
Motifs are used to construct our motif graph $G$, which forms the design basis for both our generative and predictive methods. The complexity of our grammar as conveyed by the size of the motif graph $G$ for different motif collection strategies we tried are summarized in Table 5. For the remainder of this section, we describe the Expert Annotation strategy which is our primary workflow. The strategies for obtaining motifs from literature and heuristic-based fragmentation are described in B.1 and F.1, respectively.
Grammar Complexity $(|V|,|E|)$ HOPV PTC GC
Literature N/A N/A $(96,3656)$
Expert $(329,37273)$ $(407,23145)$ N/A
Heuristic $(208,16880)$ $(279,37968)$ $(90,4095)$
Table 5: Number of nodes and edges of motif graph $G$ constructed using different annotation strategies
A.1 Expert Annotation Workflow
@@
First, experts view figures of the molecules, and indicate the bonds to break in order to segment the molecule into coherently chosen sub-fragments, shown in Figure 9. We provide a brief description of the example datasets we show here and elaborate on the rationale behind the experts’ segmentation strategy:
Table 6: Segmentation of the molecules (a) to (d) in 9. Bonds to break indicates the chemical bonds to cut to create black fragments, while the black groups and red groups listed for each molecules correspond to one another, respectively.
@@
Predictive Toxicology Challenge (PTC) (Helma et al., 2001) The small molecules are characterized by distinct functional groups known for their carcinogenic properties or liver toxicity (Miller et al., 1949; Helma et al., 2001). These groups comprise a rich variety of elements such as halides, alkylating agents, epoxides, and furan rings. (Figure 3) Therefore, we specifically segmented it into functional groups and sub-structures that contribute most to the toxicity of the compounds (Hughes et al., 2015).
@@
Defining Membership.
The Membership metric is reported in 4.2.2 after further consultation with experts, who identify the presence of Thiophene as a proxy for Membership to HOPV, and the presence of Chloride/Bromide Halides (a key indicator of toxicity) for PTC. In the case of both datasets, the Membership metric is only a sanity check that the method can produce a non-trivial number of characteristic compounds. Here’s our justification for the criteria on each dataset:
  1. 1.
    A chloroalkane (Cl-C) is the most common motif in the PTC dataset. Yet, it is still not present in a majority of structures, making the broader class of alkyl halides (Cl-C, Br-C-C) the best choice for a membership criterion. Their prevalence is attributed to their reactivity and ability to undergo metabolic activation (Leung et al., 2012), leading to the formation of highly reactive intermediates that can interact with DNA and other cellular components, potentially initiating carcinogenic processes. Although not all carcinogenic compounds will necessarily contain this class of motifs, their presence contributes a strong likelihood.
  2. 2.
    Thiophene, a 5-member ring with one sulfur group, is the most common motif in the HOPV dataset, making it the best choice for a single-motif membership criterion for HOPV. More broadly, thiophene and its derivatives are arguably the most common chemical substructure in photovoltaics due to their ability to donate electrons, resulting in particularly high highest occupied molecular orbital (HOMO) levels, along with stability, tunability of energy levels, and compatibility with film forming techniques. While not every suitable organic photovoltaic compound will contain it, the vast majority will.
In both cases, our method can easily achieve 100% membership with a slight modification to the sampling procedure: instead of iterating through every possible starting motif node, we always initialize our random walk at the membership motif. We choose not to modify our sampling procedure, and instead include this metric in Table 3 for completeness, since it is still a good sanity check for other methods to show they generate a non-trivial fraction of candidates with those motif(s).
@@ Appendix B Representing Existing Molecules as Walks on This Graph
B.1 Extracting Walks from Segmentation (HOPV, PTC)
During segmentation, we use the workflow in Appendix A.1 to segment a molecule into fragments. In doing so, we also obtain the molecule’s representation $H_{M}$ as a directed subgraph over the motif graph. The pseudocode is found in Algorithm 1.
@@ 13 Q $\leftarrow$ queue([(root,src)]);
Algorithm 2 linearizes the molecule into a directed acyclic graph (DAG). This procedure begins by finding the longest path, and choosing a consistent ordering over neighbors ($N_{G_{i}}$) to determine the random walk sequence. We elaborate on the reasoning behind this canonicalization in Appendix B.4. The DAG constraint enables our graph diffusion process to become a generator of new molecules (as will be discussed in Appendix D), in addition to capturing the distribution of existing ones. Thus, we specifically ask experts to create segmentations that are acyclic, which they naturally do in nearly all cases anyway. In the case of monomers, this canonicalization is consistent with the IUPAC nomenclature(IUPAC, 1997) of linearizing a monomer via its longest (main) chain, where $N_{G_{i}}$ should iterate over neighbor fragments that descend side chains before the consecutive fragment on the backbone of the main chain. More specifically, src and dest in Algorithm 2 correspond to the first and last group of the main chain.
B.2 Extracting Walks From Literature (Case Study of Group Contribution)
@@
The Group Contribution dataset includes a compilation of motifs characterized for gas separation, including common organic chemical functional groups as well as important scaffold functional groups such as Triptycene and its derivatives, dioxin and its derivatives, and N-methylphthalimide and PIM-1 and its derivatives (Wang et al., 2018, 2020). These functional groups contribute significantly to maintaining the structures and properties of 3D scaffold building blocks in polymer self-assembly, which in turn play a significant role in gas separation processes, i.e. the separation of $H_{2},H_{2}/N_{2},O_{2},O_{2}/N_{2},CO_{2},CO_{2}/CH_{4}$ which are common separation tasks important in gas and oil industry. The steps we take for compiling this dataset of segmented monomers are as follows:
@@
B.3 Graph Augmentation
The motif graph is the directed, multi-edge graph $G=(V,E)$. When traversing to a previously seen motif v, there is ambiguity in whether the random walk forms a cycle vs creating a copy of the previous motif and appending to the trajectory. To remove this ambiguity, the random walk traverses a duplicate node, $v_{k}$ for the latter case. A dataset of molecules and their representations, $D:=\{(M,H_{M})\}$ thus induces “an augmented version of $G$” =: $G’$ ’. For each $v\in V$, let $K_{v}=\max(\text{count}(v,H_{M})$ for M). We create duplicates for $v$ and the in/out-edges of $v$:
$V’$ ’ $\leftarrow V\cup\bigcup_{v\in V}\{v_{k}\mid k=0,\ldots,K_{v}-2\}$ (9)
$E’$ ’ $\leftarrow E\cup\bigcup_{v\in V}\{(v_{k},v’,e)\mid(v,v’,e)\in E,\forall k=0,\ldots,K_{v}-2\}$ (10)
$E’$ ’ $\leftarrow E\cup\bigcup_{v\in V}\{(v’,v_{k},e)\mid(v’,v,e)\in E,\forall k=0,\ldots,K_{v}-2\}.$ (11)
Molecule $M=(V_{M},E_{M})$ is then a rooted subgraph of $G’$ ’. In the main text, we refer to $G$ as its augmented version, to simplify the notation.
B.4 Data Augmentation
@@
As described in Algorithm 2, we linearize a molecule by first setting its “main chain” – the longest shortest path of $H_{M}$. If this happens to be part of a cycle, we disregard one edge. If there are multiple longest shortest paths, we choose the one whose first differing node comes first in our canonical ordering over the nodes of G.
@@
Appendix C Building the Motif Graph & More Related Works
Expanding on Section 3.1, we apply a subgraph-matching algorithm with pseudocode in Algorithm 3 over all pairs of motifs $v_{1},v_{2}$. This algorithm is embarrassingly parallel and runtime-efficient as the subgraph $v_{R_{l}}$ is, unless specified otherwise, a few atoms or a ring. RDKit(Landrum, 2016) provides out-of-the-box implementations for subgraph matching optimized for molecular sub-fragments like rings, enabling a significant speedup in runtime.
@@ 1 G = graph(V);
Figure 11: Following notations of main text, $\{v_{1}\}_{r_{1}}=\{1,9\},\{v_{1}\}_{r_{2}}=\{1\},\{v_{1}\}_{r_{3}}=\{9\},\{v_{1}\}_{r_{4}}=\{5\}$. $\{v_{2}\}_{r_{1}}=\{1,2,3,4,5,6\},\{v_{2}\}_{r_{2}}=\{12,13,14,15,16,17\}$. We annotate $e_{1,1}$, where $b_{1}=\{6,4,3,2,8,7\},b_{2}=\{10,7\}$ provides the “certificate” of a successful match.
@@ C.2 Connection to Graph Coarsening
Mathematically, the motif graph advocated in this work is the quotient graph of the molecular graph, under the equivalence relation defined as $u\equiv v$ if nodes $u$ and $v$ belong to the same motif. As our motifs do not overlap and jointly cover all nodes of the molecular graph, they define a partitioning of the graph. In scientific computing, collapsing each partition into a single node and retaining edges crossing partitions is called graph coarsening, which is a commonly used technique to solve large-scale problems, notably solving sparse linear systems of equations (Chen et al., 2022). Working on the coarsened version of the graph (i.e., the quotient graph) is computationally attractive as the graph size is much smaller. Moreover, when applied to machine learning problems such as graph classification, it is demonstrated that the representation learned from the quotient graph can be as predictive as that learned from the original graph (Chen et al., 2023; Ma & Chen, 2021; Cai et al., 2021). Favorably, a unique scenario of this work is that all concerned (molecular) graphs share the same set of motifs, which brings in the potential benefit of learning better molecule representations based on motif representations that form the basis of all molecules.
C.3 Connection to Random Walk Literature
Our parameterization of the random walk is by learning a graph heat diffusion process over the motif graph $G$. The relationship between graph heat diffusion and random walk has been studied before (Masuda et al., 2017), but we integrate two new ideas: 1) making the Laplacian (edge weights) learnable and dynamically adjustable, and 2) conditioning the adjustment on an order-invariant memory. The justification as to why we don’t just use autoregressive models is part of a larger discussion on the respective merits of autoregressive models vs grammar-based approaches. In data-efficient settings, previous works (Guo et al., 2023a, b) show grammar (esp. context-free grammar) work well due to the relatively small (tens/hundreds) number of examples needed to learn valid rules and derivation sequences. Meanwhile, the number of possible hidden states that autoregressive models (Li et al., 2018; You et al., 2018; Liu et al., 2018) are parameterized to learn is exponential (to the length of the sequence), and learning a good parameterization is difficult (Jin et al., 2018, 2020). We take a middle ground, combining the data-efficient advantages of context-free grammar and the expressivity of autoregressive models, by introducing a context-sensitive grammar which utilizes a set-based memory during the random walk. The set-based memory mechanism $c^{(t)}$ keeps an order-invariant memory of the nodes visited so far. Without the memory mechanism, our model becomes an order 1 Markov process. Previous literature show that higher-order random walks are required to capture temporal correlations in edge activations (Rosvall et al., 2014; Masuda et al., 2017), with a tradeoff of complexity and practicality. In the design of complex and modular structures, the order 1 Markov assumption is not sufficient (see footnote 2 in the paper). Meanwhile, higher-order models make it difficult to scale our grammar to larger motif graphs. We take a middle ground by introducing a set-based memory state, replacing the entire visit history with a summary of node visit counts. In particular, prior works study how memory mechanisms in random walks affect exploration efficiency (Fang et al., 2023; Gasieniec & Radzik, 2008) and enable negative/positive feedback (Fang et al., 2023; Pemantle, 1988). Our results in Section 4.2.2 demonstrate the efficacy of this approach.
@@ Appendix D Grammar Learning
D.1 Graph Diffusion Strategy
@@ 6       $c^{(0)}\leftarrow$ [0 for v in G];
13             $c^{(t+1)}=\frac{t}{t+1}\cdot c^{t}+\frac{1}{t+1}\cdot p^{t}$;
@@ 24             W $\leftarrow$ W - $\alpha*\frac{\text{dLoss}}{\text{dE}}$;
D.2 Visualizing Learning Process
@@
Figure 13: We show the weight evolution of the edges incidental to $G14$ on HOPV. (Left) After processing the raw dataset into random walks, we visualize the empirical distribution of edge traversals. (Middle) After learning our context-sensitive grammar, we plot the prior edge weights, i.e. the memory-free parameter $E$. (Right) We plot the transition probabilities starting at $G14$ during the random walk generation process.
In Figure 12 and Figure 13, we see our grammar’s capacity to estimate the prior edge weights, $E$, through training, as well as correct the edge weights via a memory-sensitive adjustment during the random walk. Weights of edges that are commonly traversed after $G14$ will be amplified during training, and weights of edges that visit $G14$ from another state will be diminished.
@@ Appendix E Property Prediction
E.1 Graph Neural Network Design Choices
Table 8: Hyperparameter settings for property prediction
@@ E.1 Graph Neural Network Design Choices
We apply a Graph Isomorphism Network (Xu et al., 2019) with hyperparameters in Table 8. For molecule M with representation $H_{M}$, the node-level features include: a) the Morgan fingerprint of the motif $v_{i}$ (dimension 2048), b) the memory-free weights of its out-edges (dimension $|G|$), i.e. E[i]. We also concatenate the Morgan fingerprint of M.
E.2 Bag-of-Motifs Design Choices
We obtain $|G|$-dimension motif-occurrence feature vector for each $M$. Similar to Ours, we concatenate the Morgan fingerprint of $M$ to it. We use XGBoost with $16$ estimators (boosting rounds) and maximum tree depth of $10$.
E.3 Optimization Design Choices